OpenStudy (ohmath123):

|2x+5|=4 When i solve the problem i get x=-1/2 but keeps telling me I am wrong. am i doing something wrong?

OpenStudy (isaidavila):

I think you're right because if you replace the value of x in the equation you get 4 and absoluta value of 4 is 4 Hooe it helps!

OpenStudy (anonymous):

you aer wrong because there are two answers

OpenStudy (ohmath123):

How do i get the two answers? @satellite73

OpenStudy (isaidavila):

There are two equations -(2x+5)=4 and 2x+5=4

OpenStudy (isaidavila):

No sorry is 2x+5=4 and 2x+5=-4